OpenStudy (anonymous):

Find the sum of the infinite series 4 + 3.2 + 2.56 + 2.048 +…if it exists.

OpenStudy (anonymous):

just find r

OpenStudy (anonymous):

and conclude that if r>1 no infinte sum ........ if r<1 there is

OpenStudy (anonymous):

here r = t2/t1 = 3.2/4 = 0.8 0.8<1 so there is nw using formula just find it

OpenStudy (anonymous):

@summer1nthec1ty still confused????